OpenStudy (anonymous):

find the surface area generated when rotating y=x^2 from(1,1) to (2,4) about the x axis @hartnn

hartnn (hartnn):

there's a formula for this right ?

OpenStudy (anonymous):

integral of 2pi (x or y) ds

hartnn (hartnn):

some thing like this ? \(\Large \int \limits_1^2x^2 \sqrt{1+4x^2}dx\) ?

hartnn (hartnn):

oh, i missed 2pi ...and i didn't know that...

hartnn (hartnn):

how would you find ds ?

OpenStudy (anonymous):

ds = sqrt(1+ (dy/dx))...basically the sqrt you gave in your post

hartnn (hartnn):

\(\Large 2\pi\int \limits_1^2x^2 \sqrt{1+4x^2}dx\)

OpenStudy (anonymous):

i was just a little confused about the points of integration but it's actually quite simple ( about x axis means use x coords and about y axis means use y coords right?)

hartnn (hartnn):

yes!

OpenStudy (anonymous):

should i take u=x or do this: intgrl x^2 4(1/4+x^2) and then let u=1/4+x^2?

hartnn (hartnn):

thats not so easy to integrate....

hartnn (hartnn):

since i see the form x^2 +a^2 i would have used the substitution of x = a tan t any better ideas any two of you have got ?

OpenStudy (anonymous):

i thought of letting u =1+4x^2 then du=8x =>8du=x but then the x^2 on the outside wouldn't be accounted for and i dont think theres such a thing as du^2

hartnn (hartnn):

thats correct, it won't work

OpenStudy (anonymous):

i dont follow the tan substitution you have, could you explain it out to me?

hartnn (hartnn):

when we see the form x^2+a^2 in the integral, we plug in x = a tan t so that x^2+a^2 becomes a^2 sec^2 t just a trigonometric substitution to simplify things...

OpenStudy (anonymous):

ok, so we'll be integrating x^2(1sec^2t) ?

hartnn (hartnn):

but we need to put every x as a tan t and find dx too in terms of t and tan t x= a tan t dx = .... ?

OpenStudy (anonymous):

a sec^2t

OpenStudy (anonymous):

a^2

hartnn (hartnn):

just ( a sec^2 t) dt don't forget the dt too so now whats your integral in terms of 't' ?

OpenStudy (anonymous):

is it t=a tan^-1 x?

hartnn (hartnn):

.... so, your integral becomes \(\large \int (a^2\tan^2 t)\sec t \: \: a\sec^2 t dt\) got this ?

hartnn (hartnn):

this may not be the most efficient way to integrate this, but i can't think of other way...if any o2 of you can think, then you're welcome..

hartnn (hartnn):

here a = 1/2

OpenStudy (anonymous):

i'm still a little confused, i'll revise some more trig substitution and get back to you later

hartnn (hartnn):

ok, you can ask doubts related to this right now if you have any... or you can get back later too...

OpenStudy (anonymous):

it's cool, i'll post again later if i still have questions

hartnn (hartnn):

good luck! :)